Let the PPFs of the factories look like this:

1 plant: y = 40-x

2 plant: y = 100 - 4x

3 plant: y = 60 -2x

Find the maximum number of x if you want it to be exactly the same as y.


1
Expert's answer
2021-11-28T18:05:57-0500

The set of equations are as follows

"y=50-x... equation 1"

"y=100-4x... equation 2"

"y=60-2x... equation 3"

Add the first two equations 1 and 2;

"(y=40-x )+(y=100-4x) =(2y=140-5x)"

Add equation 2 and 3;

"(y= 100-4x)+(y=60-2x) =(2y=160-6x)"

Such that you obtain two set of equations which are;

"2y=140-5x ... equation 4"

"2y=160-6x... equation 5"

Making y the subject from equation 4 you get;

"y =70-2.5x....equation 6"

Substituting for y in equation 5;

"2(70-2.5x)=160-6x"

Where you get "X=20"

Substituting for the value of x in equation 6;

"y=70-(2.5*20) =20"

"Answer = 20"

Units of x needed to produce the same number of y are 20